A man beats drum at a certain distance from a large mountain. He slowly increases the rate of beating and finds that the Echo is not heard distinctly when the drum beating is at the rate of 40 per minute. he moves by 80m toward the mountain and finds that the Echo is again not heard distinctly when the rate of the beating of the drum is 1 per second what is the initial distance of the man from the mountain? (1)120m;(2)240m;(3)270m;(4)340m

Grade:12

1 Answers

Arun
6 years ago
Please find below the solution to the asked query:

Case I. When the rate of drum beating is 40 per minute

Therefore, the interval between two drum beats = 60/40 = 1.5 s
Let the distance between the mountain and the man be x m.
Speed of sound, v=2xt=2x1.5=4x3   .....(1)
Case II. When the man moves 80 m towards the mountain

Speed of sound, v=2(x80)t=2(x80)1=2x160   .....(2)From eq. (1) and (2),2x160=4x36x480=4x2x=480x=240 m

Hope this information will clear your doubts about the numerical on echo.